Back to the NIKA2 main page

Daily Reports

Nov 22nd, Tuesday

16:00-21:00; Angel: NIKA2 cryostat checked and DAQ activated for getting selected noise traces by DJ: ok. Rainy and foggy all the shift. System check devoted to the setup for VLBI dual-receiver test (Wednesday and Thursday, 12:30-16:30, LT). No observations. Wobbler system is apparently broken (MC, SS): under study.

21:00-03:00; Jean-François: opacity > 0.8 during the first half of the run, no observation. Then it "improved" down to 0.45 in tracking the Bilal tau predictions that were promising. So we decided to have a go with nika2. [DAQ restarted at ~01:20 LT; Bad KIDs: 68, 4, 104]. However, after pointing and focus on Uranus, the opacity abruptly increased to higher than 1 till the end of the shift.

Nov 23rd, Wednesday

03:00-09:00; Longji: High opacity and fog during the first half of the shift. Opacity dropped to 0.5 and it was reasonably stable at ~07h (LT). DAQ restarted; bad KIDs: 65, 8, 90. Pointing/focus on 1226+023. El. motors current alarm was triggered several times because of the wind. Resuming at 08:45. Mean tau_225 about 0.45. Bilal takes over.

09:00-14:00; Bilal: Starting with mediocre weather (tau~0.5) and timeline variations above 50Jy! I've been looping around pointings and focus because of very poor weather, deep within clouds. At the end of the shift, moving on to calibrators to get some values. Instability going from 50Jy to 200Jy, easing down to 10Jy at the end of the shift.

15:00-21:00; Angel: tau_225 < 0.3 over the shift. The atmosphere was quite unstable till ~15:30 (p2p variations oscillated between +/-5 and 20 Jy/beam at 2 mm; slightly enhanced beam: e.g. FHWM 18.6 x 19.5 sq.-arcsec). After some stabilization, we got a calib_1scan on NGC7027 and CRL2688 (scans 202211s127 and s129), and a skydip (s132). Switching to project 160-16 (IMEGIN). We stopped for over 20 min at 16:50 (LT) for checking the Wobbler system (SS). We took advantage of that test for fresh resonances (DAQ restarted; bad KIDs: 52, 7, 73) under more stable conditions and refocusing. At last, we observed NGC7331 for ~1.5 h, distributed in 4.5 reps., until the target reached high elevations. Then, we moved to NGC337 (IMEGIN) and observed this galaxy for 1.2 h (8 reps.). At 20:40 we pointed to Neptune (under nikas-22) and launched a calib_1scan (s180).

21:00-03:00; Aurore: Calibrations (beammap [20221124s101], calib_1scan and skydip) were performed on Uranus at the beginning of the shift. Then I switched to GASTON for 2 scans, however even though the atmosphere was stable at the beginning, it deteriorated during the scans, so we will probably have to throw them away. After the stability improved, I relaunched another scan but the atmosphere deteriorated again. I redid a calibration round (calib_1scan and skydip) on Uranus after that, since the sources of GASTON and 008-16 were too high and none of the sources of 089-22 were up. Unfortunately I missed the calibrations on CRL618, as it was also too high. Towards the end of the shift, I tried to go to the one source of 089-22 that was up, hoping that by the time the telescope reached the source, the stability would have improved, but in fact, it was worse. Together with the telescope being stuck in elevation about 4-5 times, it was quite a frustrating night, despite a reasonable opacity (around 0.3 most of the time, but with frequent spikes at ~ 0.4-0.5).

Nov 24th, Thursday

03:00-09:00; Longji: The observation started with 2 reps on IC348-1 under project 008-22. After that, calib_1scan on CRL618 and ALFORI, as well as a skydip were made. The conditions were not stable enough for N2CLS so I switched to 089-22 to observe AR2747 for 1 rep. [resonances worked at 07:30; bad KIDs: 46, 6, 65]. The last few hours are spent on COSMOS observations under N2CLS project (192-16), with improved stability.

09:00-12:30; Bilal: Starting where Longji left, with observations of COSMOS in good weather conditions. tau_225~0.25 and timeline variations around 2-3Jy. A skydip got at 10:00 (20221124s70). Four stuck scan canceled (20221124s65, 66, 67 and 68). I continued with LPSZ observations (PSZ2G086.9) for two repetitions. After a reasonable pointing and focus session, we move to IMEGIN, to observe NGC4254 for 3 repetitions, sharing the good weather with everyone.

12:30-16:30; Pablo: EMIR dual-receiver FPT tests for VLBI.

16:30-21:55; Angel: DAQ restarted at ~16:30 (bad KIDs 45, 8, 60). Opacity < 0.2 at the beginning of the shift. Quite stable atmosphere (p2p variations below +/-4 to 5 Jy/beam at 2 mm). Cluster PSZ2G045 was observed for 199-16 (requested reps. completed). Switching to nikas-22: a calib_1scan on Neptune (20221124s176) + a skydip (s177). We changed to project 160-16 (IMEGIN). After pointing/focus and the first try on NGC337 (07:10), a tsubreflector-trace alarm triggered. Fixed by PM and SA. After a second try on NGC337 (~07:30), synchronization problems were noticed thanks to the plots of PIIC monitors. DAQ main window has disappeared, despite the remaining systems were running, but DAQ's VAntenna was turned on red, and communications with mrt-lx1 had apparently failed. DAQ restarted without unusual warning messages at 20:45 (bad KIDs 46, 6, 66) after several tries. DJ was informed and checked. Observations resumed at ~21:00. We lost about 1.5 h. Scans probably compromised: 188-189 and 193-194 (NGC337; 19 min in total). After pointing/focus on Uranus, a calib_1scan on Uranus (s205) was used for further tests in all PIIC monitors. No synchronization problems anymore. Quite a stable atmosphere; p2p variations of +/-1 Jy/beam; <rms> close to 45 mJy/beam (2 mm). A beam map (s207) and a skydip (s208) were obtained after that.

22:00-03:00; Karine: Tau was very good ~0.25 and stable throughout the shift. I started on project 122-16 (GASTON) with a pointing and focus on Uranus and 1 rep on L1521e. The stability conditions, which were OK at first have then deteriorated along the shift. During the two scans on L1521e, the peak-to-peak variations increased from ~ 10/20 mJy/beam up to more than 20 mJy/beam and the rms was > 156 mJy/beam. After several pointings and a focus on Uranus I did a calib_1scan on Uranus and another one on ALFORI. I switched to project 089-22 and did a rep on J1104_5040 (p2p variations +/- 10-20 mJy/beam, rms ~200 mJy/beam,tau ~0.25). Stability conditions deteriorated to +/- 50 mJy at ~2 am. Pointing on Uranus and focus (which was good). We then switched to EMIR at the very end of the shift.

Nov 25th, Friday

03:00-09:00; Longji, Angel (till 06:30): Tau about 0.25 at the beginning of the shift. Preparing setup for observing SU Aur of EMIR backup project 062-22. Tuning (E0+E1) ok, <pwv> ~2 mm in the l.o.s. (MRTCAL/FTS). Unable to get data for MIRA or MRTCAL plots using BBC (and even NBC) backend. Trying to fix that. Opacity increasing in the meantime. High altitude clouds. Updated at 06:05 (LT).

09:00-15:00; Laurence, Angel: Problem impacting the wobbler switching system for EMIR observation, at the beginning of the shift. Solved by Angel and the IRAM staff. The opacity was around 0.4 and decreasing. Starting observation using EMIR around 10:30 (backup program 083-22, 3 repetitions on source 29349-high). The opacity continues to slowly improve. When tau_225 was oscillating around 0.25, we decided to switch back to NIKA2. Atmospheric instabilities prevented us from tuning the detectors. Angel did multiple sweeps trying pointing toward the clearest part of the sky. Eventually, we got reasonable resonance measurements, but 130 "bad KIDs" in A1, 100 bad KIDs in A3, and 5 in A2. We continued trying until we got convinced that this is the best we could do given the atmospheric conditions. NIKA2 observations can begin! However, another surprise awaited us after the first pointing. We chose a very bright QSO but got a flat field in PIIC! We tried a couple of other sources, including sources at higher elevations, to no avail. After several double-cross-checks, we concluded the issue was antenna-related, most likely the secondary mirror position. Ignacio intervened to check the antenna system. The NFS was restarted, and the wobbler blockage was double-checked. Eventually, we got a correct pointing around 13:00 LT. We used the rest of the shift to perform calib_1scan on secondary calibrators.

15:00-21:00; Angel: Opacity < 0.12 over the whole shift, except for an isolated spike, very stable atmosphere (p2p variations between +/-1 to 6 Jy/beam at 2 mm). Mean rms was steadily below 45 mJy/beam, sometimes about 38 mJy/beam in that band. TAPAS records stopped today at ~2 a.m.: reported to the IT group; the recovering process is ongoing with the appropriate priority in order to avoid mrt-lx3 overloading (e.g. PIIC-QL is running there). Taking advantage of the stability, new resonances were obtained at ~15:50; bad KIDS: 25 (ar1), 6 (ar2), 47 (ar3). NGC7331 of IMEGIN was observed for 2.23 h in total (6 reps with 60 and 45 arcsec/s scan speeds). Then we switched to project 199-16 (LPSZ): two reps. on PSZ2G045. Switching to nikas-22 account. A calib_1scan was launched on Neptune (20221125s164) + a skydip (s165). Swapping to Uranus for a calib_1scan (s173) and updating the inclinometers after pointing and focus. Switching to project IMEGIN. Target NGC337 was observed twice for 27 min.

21:00-03:00; Hélène: The conditions were excellent and the shift ran smoothly, with tau ranging from 0.1 to 0.13 most of the time, with a short excursion up to 0.2, and a very stable atmosphere. The raw science data at 2mm had an rms between 2 and 3 Jy/beam most often, 5 Jy/beam at maximum. The processed data had a rms around 40 mJy/beam. The focus changed quite a lot during the night, and was checked relatively often. Calibrations done: a beammap on Uranus (20221125s210), 2 calib_1scan on Uranus (20221126s1 and 20221126s12), a calib_1scan on Alfori (20221126s5) and a skydip. Science observations: 2 scans on N337 to continue the program of the previous shift, and 10 scans on L1521E (two of them at 50"/s, the rest at 60"/s). Note on the RTA: it was impossible to obtain a good fit with the IDL RTA for a pointing done on Alfori, while the PIIC RTA gave good results for the same scan. The RTA takes a very long time after a scan is processed before giving back the prompt, which is a drawback for focus sequences.

Nov 26th, Saturday

03:00-09:00; Longji: Observation started from 2 reps on PSZ2G133 under project 199-16. Then I switched to COSMOS (192-16) during the rest time of the shift. New focus scans are made at 4:20 and 6:50. The opacity was stable at around 0.1 throughout the shift. RMSs of 5 Jy/beam in the raw data at 2mm at maximum also indicates good stability.

09:00-15:00; Juan: new resonances at 09h; bad KIDS: 25 (ar1), 6 (ar2), 47 (ar3). Opacity < 0.13 over the shift, quite stable atmosphere. On past noon, the beam sizes broadened enough to continue on mainly source-detection projects. About science scans: continuing on COSMOS field (1.5 reps. = 3 scans) for project 192-16, 4 reps. on PSZ2-G086.9 (project 199-16), and 1 rep. on J1327+4846 (project 089-22). As for calibrations: a calib_1scan on NGC7027 (20221127s119) and a skydip (s120).

15:00-21:00; Angel: Opacity ~0.13 at the beginning of the shift. Recovering reasonable beam sizes after checking with focus/pointing on 1633+382 (under nikas-22). Highly stable (p2p +/- 1 Jy/b at 2 mm); beam sizes: 17.2 x 17.9 sq.-arcsec (2 mm); 11.0 x 11.6 sq.-arcsec (1 mm). Switching to project 199-16: 2 reps. on PSZ2-G086.9 (required observations complete; please check). Then we observed NGC7331 of project 160-16 for 1h (4.5 reps.) up to max. elevation. After that, we obtained a calib_1scan (20221126s182) on Neptune and a skydip (s183). Coming back to project 160-16. Three reps. (~45 min) on NGC337. Unstable focus in the past ~1 hour. Switching to nikas-22 for monitoring. Hélène takes over.

21:00-03:00; Hélène: The conditions were pretty much the same as yesterday, with opacities around 0.1 most of the time (with short excursions up to 0.2 or 0.3), rising very slowly toward the end of the shift, rms in raw 2mm science data between 0.9 and 3 Jy/beam mainly, maximum 5.3 Jy/beam, and rms in processed 2mm science data around 40 mJy/beam. The focus was much more stable than last night. Calibrations done: two beam maps on Uranus (20221126s244 and 20221127s9), a calib_1scan on Uranus (20221127s1), and a skydip. Science observations: 2 repetitions (8 scans) on ACTJ0119 (project 199-16), and 5 repetitions (10 scans) on L1521E (4 scans at 45"/s, 6 scans at 60"/s).

Nov 27th, Sunday

03:00-09:00; Longji: Observation started with the continuation of the observations of the previous shift for 0.5h. After focus checkon 0316+413, the cluster PSZ2G133 (199-16) were observed for 3 reps. A new focus was made again on 1226+023. Then COSMOS field is observed for 6 reps (~3 hours), with an additional focus on 1226+023 in the middle. New resonances at ~08h; bad KIDS: 33 (ar1), 6 (ar2), 57 (ar3). During the shift, the opacity was stable with values at around 0.15.

09:00-14:30; Angel: Opacity about 0.15 at the beginning of the shift, pretty stable (p2p < +/- 5 Jy/b and <rms> < 42 mJy/b (2 mm) over the whole shift), light cirrus gradually covering the sky. Starting by targetting NGC4254 (scan speed: 60arcsec/s), project 160-16. Highly stable; p2p < +/- 1 Jy/b and <rms> < 40 mJy/b (2 mm); FWHM 18.2" (2 mm), 11.7" according to the IDL RTA, despite good focus. Checking beam widths after 2 reps. on NGC4254. Switching to nikas-22 for testing. Focus is stable. After focusing on 1226+023, we got FWHM_PIIC of 17.2 x 19.0 sq.-arcsec (2 mm), 10.6 x 12.3 sq.-arcsec (1 mm), and FWHM_IDL: 18.4 arcsec (2 mm), 12.1 arcsec (1 mm), so a bit broadened. However, after pointing on the neighbor QSO 1253+055, we got FWHM_PIIC: 17.6 x 18.4 sq.-arcsec (2 mm), 11.0 x 12.3 sq.-arcsec (1 mm), and FWHM_IDL: 17.7 arcsec (2 mm), 11.8 arcsec (1 mm). A source effect is evident. In any case, beam widths are no ideal. We further check the evolution of temperatures of the back-structure and everything looks as expected, compared with the temperature evolution of the past days. Environmental temperature is about 5C and increasing, with no wind, despite being hazy. Telescope has been pointing until now at 90+ deg. only in azimuth from the position of the Sun. We continued on NGC4254 by adding 3 reps. more. Beam widths checked on 1253+085 under nikas-22: FWHM_PIIC: 17.1 x 17.2 sq.-arcsec (2 mm), 10.9 x 11.8 sq.-arcsec (1 mm); FWHM_IDL: 17.7 arcsec (2 mm), 11.8 arcsec (1 mm). Pretty nice. Continuing on NGC4254 science scans after pointing at 1226+023: 1.5 reps. Total time on NGC4254 during this shift: 2.2 h (6.5 reps.). Switching to project 199-16 for at least 1 rep. on PSZ2G133. Pretty focus and pointing on 1253-055. Quite stable, but we got FWHM_PIIC of 17.2 x 18.4 sq.-arcsec (2 mm), 11.3 x 13.7 sq.-arcsec (1 mm), and really the provided repetition would end below El.=40 deg. Canceling the attempt. Taking advantage of the last focus/pointing obtained to tackle a project mainly addressed to the detection of compact sources: 095-22. Got 9 Az. scans on AR2747. Finishing with a skydip (202211s119). Starting setup for project POLAMI.

14:30-16:30: Juan Escudero for POLAMI.

16:30-21:00; Juan: Restarting DAQ at 16:50; bad KIDS: 114 (ar1), 4 (ar2), 113 (ar3). Quite difficult to get better resonance sets, unstable atmosphere. Unable to get reference pixels. We retried at 17:30: bad KIDS: 41 (ar1), 7 (ar2), 63 (ar3). Pointing corrections ok. Switching to project 160-16. One effective rep. on NGC7331. Too high for continuing on this target. Switching to nikas-22. Pointing/focus, and a calib_1scan on Neptune [20221127s167] + a skydip [s168]. Switching to project 160-16 again: 4 reps. on NGC337 (first two with a bit unstable atmosphere). At the end of the shift, I got a calib_1scan on MWC349 [20221127s190] and CRL2688 [s192] + a skydip [s193].

21:00-03:00; Aurore: Conditions were good on average, with opacities mostly between 0.2 and 0.3. Stability could have been better but was acceptable. I did calib_1scan on CRL618, ALFORI, and a whole round of calibrations on Uranus (skydip [20221127s241], a calib_1scan [20221128s1], and a beam map [s8]) between around 00:00 - 02:00 when target sources were either too high or too low. Focus was quite stable at the beginning but changed suddenly around midnight (I repeated it a few times to be sure it was correct and the results were consistent). Sciencewise, I did a repetition on IC348-1 (008-22), one scan on L1521E (122-16), one repetition on B211 (122-16) and one repetition on B211-213. I still need to fill in the README files to keep track.

Nov 28th, Monday

03:00-09:00; Longji: The observation started with two series of focus scans on Uranus and 0316+413, with opacity changing significantly. However, it was difficult to find a consistent solution and proper beam size. Other sources, including Mars, 0851+202, and 1055+018 were used to check the focus in the following hours, with solutions consistently between -0.4 and -0.5. During this time, the beam at 2mm was barely acceptable with a size of ~17.5"x19", while the 1mm beam was significantly elongated, with a size of 11"x15" even at the best solution. This prevents us to observe PSZ2G133 (199-16) and COSMOS (192-16) as originally planned. New resonances at 05:40; bad KIDS: 26 (ar1), 6 (ar2), 43 (ar3). At 7 am we decided to switch to 089-22 mostly aiming at source detection and observing J1030+0135 with the highest priority. At 8am we check the focus on 1253-055 and find an improvement but still elongated beam. We finally switched to COSMOS at 8:30am and observe for 2 reps (1 hour).

09:00-14:30; Angel: Opacity about 0.2 at the beginning of the shift. Last rep. on COSMOS running. Switching to nikas-22 after that for testing beam sizes, and previous focus correction. Nicely stable; from pointing on 1253-055 and after focus (nikas-22), FWHM_PIIC 17.6 x 18.2 sq.-arcsec (2 mm), 11.8 x 12.5 sq.-arcsec (1 mm); FWHM_IDL 17.7 arcsec (2 mm), 11.9 arcsec (1 mm). Switching to project 199-19. Nice beams ~1 h after the last focus correction: FWHM_IDL 17.5 arcsec (2 mm), 11.6 arcsec (1 mm). At 11:03 (LT) a science scan got stuck. The telescope does not respond to the usual commands after canceling scans. IT group and IR were contacted. Resuming observations at 11:48 after restarting VAC1, at last. Finally, we obtained 5 reps. on PSZG133 (requested reps. done). Pointing correction from 1156+295, scan s171, under nikas-22: Quite stable. FWHM_PIIC 17.3 x 17.7 sq.-arcsec (2 mm), 11.2 x 11.5 sq.-arcsec (1 mm); FWHM_IDL 17.5 arcsec (2 mm), 11.5 arcsec (1 mm). Switching to project 160-16, 1 rep. NGC4594 (scans s172 and s173). The scan s172 is apparently fine (according to IDL-RTA), but the last subscan appears incomplete in the IDL plots. But starting from scan s173 to s176, something was wrong according to DAQ monitors: Error MJD antenna. DAQ partially crashed. The creation of imbFITS stopped after scan s171. The plan to get calibrations (MWC349 and CRL2688) + a skydip was aborted. DAQ was forced to close. Preparing EMIR for POLAMI session.

14:30-16:30: Juan Escudero for POLAMI.

16:30-21:00; Angel: Beam widths are a bit broadened. Refocusing. Quite stable. Opacity below 0.2 (with occasional spikes) over the shift. Pointing/focus on 2234+282. FWHM 17.2 x 18.2" (2 mm). Switching to project 160-16. Started 1 rep. (60 arcsec/s) on NGC7331: telescope's NCS error when sub-scan 29 of 20221128s229 was running (18:00). Canceled. The target is too high to continue after fixing the problem. Switching to nikas-22; point/focus on Neptune. FWHM_PIIC 17.1 x 18.4 sq.-arcsec (2 mm), 11.7 x 12.5 sq.-arcsec (1 mm). Telescope stopped at 18:40 for rebooting VAC1 (asked by IR). Restarting DAQ at 19:10; bad KIDS: 25 (ar1), 6 (ar2), 51 (ar3). Waiting for a while for restarting Antenna Mount Drive services. It was needed to restart DAQ again at 19:30; bad KIDS: 25 (ar1), 6 (ar2), 51 (ar3). Tracking to K3-50A for updating inclinometers. Resuming the observations at 19:50: about 1h50m lost. Pointing/focus on Neptune + calib_1scan [20221128s252] + skydip [s253]. Pointing/focus on MWC349. FWHM 17.5" (2 mm) and 12.8" (1 mm). A series of calib_1scan on MWC349, NGC7027, and CRL2688 [s262-s264] were obtained after that.

21:00-03:00; Jean-François: Fair atmospheric conditions (tau~0.2, peak-to-peak fluctuations < 10 Jy/b, a few huge spikes that must not be atmospheric). Project 122-16: B213, B111 at 40"/s, B211_13, one repeat each. Project 199-16: ACTJ0119, 4 repeats while above 40deg of elevation as requested. Calibrations: beam map on Uranus [20221129s1], a skydip [s2], and a calib_1scan on ALFORI [s4]. Focus was checked twice (once just before the beam map) and was very stable at -0.58. Beam widths were standard at 17.5" x 17.5" at 2mm, and slightly broader than standard at 1mm (12" x 13.5").

Nov 29th, Tuesday

03:00-09:00; Longji: Continued the observation of Jean-François for 30 mins. A new focus was made on 0316+413 afterward, with a slightly elongated beam at 1mm. After that, we switched to PSZG1226 (199-16) and observed for 2 reps. At 5:10 we started the observation on COSMOS (192-16). The opacity was around 0.25 and gradually increases to ~0.3 with time but the stability was excellent, with peak-to-peak variation in pointing scans of less than +-2 Jy/beam at 2mm. At 8:30 we started the observations of AK4-224 of project 091-22.

09:00-10:30; Angel, Miguel: Opacity about 0.3, with a large spike that apparently did not affect the involved scan. Observations of AK4-224 continued. Despite the Sun being about 60 deg. from that target, the focus was stable and beam sizes were around the standards. Baseline p2p variations smaller than +/-5 Jy/beam and <rms> < 55 mJy/beam (2 mm) for all scans. Finally, we got 1.75 reps. on AK4-224 under quite stable conditions.

DailyReportsNika2Pool221122 (last edited 2022-11-30 12:11:46 by NikaBolometer)